Two forces, F⃗ 1 and F⃗ 2, act at a point. F⃗ 1 has a magnitude of 9.60 N and is directed at an angle of 59.0 ∘ above the negative x axis in the second quadrant. F⃗ 2 has a magnitude of 5.40 N and is directed at an angle of 53.6 ∘ below the negative x axis in the third quadrant.

What is the x component of the resultant force? Express your answer in newtons.

What is the y component of the resultant force? Express your answer in newtons.

What is the magnitude of the resultant force? Express your answer in newtons.
Answer #1

here,

F1 = 9.6 N * ( - cos(59) i + sin(59) j)

F1 = - 4.94 i N + 8.23 j N

F2 = 5.4 N * ( - cos(53.6) i - sin(53.6) j)

F2 = - 3.2 i N - 4.35 j N

the resulting force F = F1 + F2

F = - 8.14 i N + 3.88 j N

the x-component of the resulting force is (-8.14 N)

the y-component of the resulting force is (3.88 N)

the magnitude of resulting force , |F| = sqrt(Fx^2 + Fy^2 )

|F| = sqrt(8.14^2 + 3.88^2) = 9.02 N